OpenStudy (anonymous):

Solve for y: 2 over 3y + 2 = 8 y = 11 y = -11 y = -9 y = 9

OpenStudy (anonymous):

y=9

OpenStudy (anonymous):

Wait are you saying 2/(3y) + 2 or 2/(3y+2)?

OpenStudy (anonymous):

\[\frac{2}{3y+2}=8\] multiply both sides by \(3y+2\) to clear the fraction. use parentheses then multiply out using the distributive law

OpenStudy (anonymous):

I did it as 2/(3y) +2

OpenStudy (anonymous):

2/3y+2=8

OpenStudy (anonymous):

oh maybe i am wrong

OpenStudy (anonymous):

I'm sorry about that.

OpenStudy (anonymous):

Ok then my answer was right. Satellite would have been right otherwise

OpenStudy (anonymous):

Actually I'm still confused about what the initial question was. Either way, you have the right answer

OpenStudy (anonymous):

now i am confused

OpenStudy (anonymous):

Here I will re write it.

OpenStudy (anonymous):

i bet it is \(\frac{2}{3}y+2=8\) am i right?

OpenStudy (anonymous):

] Solve for y: 2/3y + 2 = 8

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Haha ok then

OpenStudy (anonymous):

So what's the answer? I'm lost.

OpenStudy (anonymous):

\[\frac{2}{3}x+2=8\] first step, subtract 2

OpenStudy (anonymous):

you get \[\frac{2}{3}x=6\]

OpenStudy (anonymous):

My answer of y=9 is right.

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

then multiply both sides by the reciprocal of \(\frac{2}{3}\) which is \(\frac{3}{2}\) that will give you \(1x\) on the left, or as we now know \(x\)

OpenStudy (anonymous):

\[\frac{3}{2}\times \frac{2}{3}x=6\times \frac{3}{2}\] \[x=6\times \frac{3}{2}=3\times 3=9\] as elly said

OpenStudy (anonymous):

Thanks (: